OpenStudy (anonymous):

can someone help me why sin 45 + cos 45 = √2?

OpenStudy (campbell_st):

well using an exact ratio \[\sin(45) = \cos(45) = \frac{1}{\sqrt{2}}\] so then \[\sin(45) + \cos(45) = \frac{1}{\sqrt{2}} + \frac{1}{\sqrt{2}} = \frac{2}{\sqrt{2}}\] the denominator has been rationalised.. \[\frac{2}{\sqrt{2}} \times \frac{\sqrt{2}}{\sqrt{2}} = \frac{2\sqrt{2}}{2}\] hope this makes sense.

OpenStudy (anonymous):

okay iwas staring at this for 3 minutes but i still didn't get it >< im so sorry. isnt cos and sin 45 = √2/2 ?

OpenStudy (campbell_st):

nope its \[\frac{2}{\sqrt{2}}\]

OpenStudy (campbell_st):

from the exact value triangles